OpenStudy (anonymous):

Bacteria can divide every 20 minutes, so 1 bacterium can multiply to 2 in 20 minutes, 4 in 40 minutes, and so on. How man bacteria will there be in 6 hours? Write your answer using exponents, and then evaluate. I only want help not the answer.

OpenStudy (anonymous):

2^9

OpenStudy (anonymous):

512

OpenStudy (anonymous):

sorry didn't see the last part

OpenStudy (anonymous):

yeah well according to the back of the textbook it's 2^18

OpenStudy (anonymous):

stupid textbook

OpenStudy (anonymous):

ohhh... 6 hours... I did 3 hours

OpenStudy (anonymous):

lol that explains it

OpenStudy (anonymous):

But how did you find the answer that's what I'm after

OpenStudy (anonymous):

okay, so the key is to look at everything in terms of 2

OpenStudy (anonymous):

ok, then?

OpenStudy (anonymous):

at first: when the time is 0 the number of bacteria are 1 when the time is 20 the number of bacteria is 2 when the time is 40 the number of bacteria is 4 2^0=1 2^1=2 2^2=4

OpenStudy (anonymous):

see the pattern? Take the time... divide it by 20... use that number, and raise two to that number. there are 360 minutes in 6 hours. 360 minutes/20 minutes =18 2^18 is the answer

OpenStudy (anonymous):

but how is 2^0=1 2 = 2 not 1

OpenStudy (anonymous):

whenever a number is raised to the power of 0 it equals 1 whenever it is raised to the first power it equals itself.

OpenStudy (anonymous):

Oh ok, the teacher didn't tell us that :P